If Leah is 6 years older than her sister, Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?
Answer by AnlytcPhil(1116)


If Leah is 6 years older than her sister, Sue, and John
is 5 years older than Leah, and the total of their ages is
41. Then how old is Sue?
`
>>...Leah is 6 years older than...Sue,...<<

Translation: L = S + 6

>>...John is 5 years older than Leah...<<

Translation: J = L + 5

>>...and the total of their ages is 41...<<

Translation L + S + J = 41
`
So you have the system:
`
L = S + 6
J = L + 5
L + S + J = 41
`
Can you solve this syetem? If not post again:
`
L - S ` ` = 6
-L ` ` + J = 5
L + S + J = 41
`
Leah's age = L = 14, Sue's age = S = 8, John's age = J = 19
`
Edwin